In an alcohol-in-glass thermometer, the alcohol column has length 12.55 cm at 0.0 ∘C and length 22.96 cm at 100.0 ∘C.

Part A: What is the temperature if the column has length 18.64 cm?

Part B: What is the temperature if the column has length 14.21 cm?

Please answer both in degrees Celsius.
Answer #1

SOLUTION :

A)

The change of length is 22.96 - 12.55 = 10.41 cm,

So the change per degree is 10.41 / 100 = 0.1041 cm/deg.

Change in length = 18.64 - 12.55 = 6.09

Temperature = 6.09/0.1041 = 58.5°C

B)

Here change in length is 14.21 - 12.55 = 1.66

Temperature = 1.66 / 0.1041 = 15.94°C